The density of a solid decreases by 0.5percent  when heated from 0 degree Celsius to 100 degree Celsius. What is coefficient of linear expansion of solid?

Grade:11

1 Answers

Eshan
5 years ago
\dfrac{d_f}{d_i}=0.995
\implies \dfrac{V_i}{V_f}=0.995
V_f=V_i(1+\gamma\Delta T)
\implies \gamma=5\times 10^{-5}/^{\circ}C=3\alpha
\implies \alpha=\dfrac{5}{3}\times 10^{-5}/\circ C